# Reconciling aging and slip state evolutions from laboratory-derived canons of rate-and-state frictionF

### v2

@@ 4.5 Incorporation of the strength minimum
This state variable $\tilde{\theta}$ always has the same $\Phi$ dependency (eq. 16) as that of $\theta$ limited to the range $\theta\gg\theta_{\rm X}$ (eq. 12) mentioned above. Therefore, even for $\theta\ll\theta_{\rm X}$, the value of $\tilde{\theta}$ can be well constrained, and hence, we are able to derive meaningful evolution laws for $\tilde{\theta}$. In fact, though a bit counter-intuitive, varying rate of $\theta$ (i.e., the evolution law in $\theta$ notation) is as well constrained as that of $\tilde{\theta}$ through $\dot{\theta}=\dot{\tilde{\theta}}$, despite the weak constraint on the value of $\theta$ for $\theta\ll\theta_{\rm X}$.
Now, we constrain the evolution law according to the canonical behaviors, this time considering the behaviors in the range $\theta\lesssim\theta_{\rm X}\cup\theta_{\rm SS}\lesssim\theta_{\rm X}$ as well. We begin with the SS canon, as we did for the development for $\theta,\theta_{\rm SS}\gg\theta_{\rm X}$. For the entire range of $V$, including $V\gg L/\theta_{\rm X}$, $V\sim L/\theta_{\rm X}$, and $V\ll L/\theta_{\rm X}$, Weeks (1993) have shown that their experimental data of the steady-state shear stress $\tau_{\rm SS}$ are well fitted by eq. (10), namely the direct-effect term plus steady-state strength, if the steady-state strength has the following velocity dependence involving a high-speed cutoff:
@@
This satisfies both the experimental requirement (eq. 26) for the SS test when $\theta_{\rm SS}\gg\theta_{\rm X}$ and the strength minimum (eq. 14). Through eq. (15), eq. (106) is equivalent to $\Omega_{\rm SS}=1$ (eq. 31). Therefore, considering the above-mentioned result of Weeks (1993), we regard eq. (31) as the SS canon for the entire range of $\theta$ and $\theta_{\rm SS}$.
The VS requirement (eq. 43) holds for any $\Phi$ and $\Phi_{\rm SS}$. Thus, no modification is necessary to incorporate the strength minimum. This VS requirement (eq. 43) is the only constraint on $f_{\delta}$, except that it must be negligible at the NSC, which is automatically met as long as eqs. (43) on $f_{\delta}$ and (65) on $f_{t|{\rm NSC}}$ hold. Hence, we can adopt the associated slip-dependent evolution function $f_{\delta{\rm|VS}}$ of eq. (43), which involves the effect of $\theta_{\rm X}$, as the evolution law $f_{\delta}$ satisfying the VS canon (eq. 38) for the entire range of $\theta$ and $\theta_{\rm SS}$. For convenience in subsequent derivations, we introduce the following variable:
@@
Also, purely time-dependent healing under the NSC has been already constrained by eq. (65), which already considers the effect of $\theta_{\rm X}$ and is therefore valid for any $\theta$ and $\theta_{\rm SS}$:
@@
Thus, the only missing is $f_{t}$ under the non-NSC. Similar to our derivation of eq. (83) for $\theta,\theta_{\rm SS}\gg\theta_{\rm X}$, the asymptotic form of $f_{t}$ under the non-NSC is constrained by the VS constraint of eq. (44). Using eq. (112), eq. (44) can be rewritten into
@@
When we derive eq. (112) from eq. (43), we have removed the conditional subscript ||||VS on the left-hand side of eq. (112) so as to indicate that eq. (112) represents the $f_{\delta}$ part of the evolution law being developed, rather than the description of constraints in a specific type of experiments. Unlike eq. (112), eq. (113) is still an experimental constraint on the evolution law, so we keep ||||VS on the left-hand side of eq. (113). This is for the pragmatic reason of distinguishing it from another constraint but for the NSC (eq. 65).
@@
Thus, regarding the constraint on $f_{t}$, the NSC ($|f_{t}|\gg|f_{\delta}V|$) gives the range where the SHS canon applies while eq. (44) ($|f_{t}|\ll|f_{\delta}V|$) gives the range where the VS canon applies. On the other hand, regarding $f_{\delta}$, since the definition of the NSC ($|f_{t}|\gg|f_{\delta}V|$) means that $f_{\delta}$ is negligible, the explicit form of $f_{\delta}$ under the NSC is poorly constrained [from $\mathcal{O}(B)$ variations in $\Phi$]. However, this conversely means that we can assume eq. (112) for the entire range including the NSC, without causing any relevant differences in the resultant behavior of $\Phi$. For this reason, in the present derivation, we can apply eq. (112) for the entire range. Similarly, noting that $f_{t}$ does not affect the evolution as long as $|f_{t}|\ll|f_{\delta}V|$ (eq. 44) holds, we can modify the constraint of eq. (113) to the following stricter one:
$f_{t|{\rm VS}}=o[(L/D_{\rm c})(\tilde{\Omega}/\tilde{\Omega}_{\rm SS})\ln(\tilde{\Omega}/\tilde{\Omega}_{\rm SS})].$ (114)
@@
where $f_{t}$ is more strongly muted than in eq. (113) because $\tilde{\Omega}_{\rm SS}\geq\Omega_{\rm SS}=1$ as seen from eqs. (31) and (111). Note eq. (114) is a sufficient condition of eq. (113). The use of eq. (114) facilitates subsequent derivation.
@@
Using eq. (114), we can derive the evolution law by a logic parallel to that in §4.2. Specifically, by noting that $\tilde{\Omega}/\tilde{\Omega}_{\rm SS}$ connects to $\Omega$ when $\theta,\theta_{\rm SS}\gg\theta_{\rm X}$ (eqs. 105 and 111), we can regard eq. (114) as a generalized form of $f_{t|{\rm VS}}=o[(L/D_{\rm c})\Omega\ln\Omega]$ derived from eqs. (44) and (72), where $\Omega$ (i.e., $\Omega/\Omega_{\rm SS}$ given eq. 31) has been replaced by $\tilde{\Omega}/\tilde{\Omega}_{\rm SS}$. Therefore, $f_{t}$ satisfying eqs. (65) and (114) can be derived by repeating the same procedure as in §4.2, with $\Omega$ replaced by $\tilde{\Omega}/\tilde{\Omega}_{\rm SS}$.
@@
Since the right-hand side of eq. (114) is a unary function of $\tilde{\Omega}/\tilde{\Omega}_{\rm SS}$, eq. (114) is equivalent to that $f_{t}$ equals the right-hand side of eq. (114) in the range of $\tilde{\Omega}/\tilde{\Omega}_{\rm SS}$ that can occur in VS tests. In ideal VS tests, where the interface slip rate instantly changes from $V_{\rm before}$ to $V_{\rm after}$, the farthest deviation of the strength from the steady state occurs immediately after the step, and hence we see that $|\ln(\tilde{\Omega}/\tilde{\Omega}_{\rm SS})|\leq|\ln[(L/V_{\rm before}+\theta_{\rm X})/(L/V_{\rm after}+\theta_{\rm X})|$, using eqs. (105) and (110). This inequality obtained for ideal VS tests is approximately valid for realistic (non-ideal) VS tests. Hence, eq. (114) can be rewritten into
@@
$f_{t}=o\left[\frac{L}{D_{\rm c}}\frac{\tilde{\Omega}}{\tilde{\Omega}_{\rm SS}}\ln\frac{\tilde{\Omega}}{\tilde{\Omega}_{\rm SS}}\right]\hskip 10.0pt\left(\left(\ln\frac{\min(V_{\rm after},V_{\rm X})}{\min(V_{\rm before},V_{\rm X})}\right)\left/\left(\ln\frac{\tilde{\Omega}}{\tilde{\Omega}_{\rm SS}}\right)\gtrsim\right.1\right)$ (115)
@@
Lastly, we look for $f_{t}$ satisfying both eq. (65) (constraint for the NSC) and eq. (115) (constraint for $f_{t}\ll|f_{\delta}V|$). Since the right-hand side of eq. (115) is a unary function of $\tilde{\Omega}/\tilde{\Omega}_{\rm SS}$, we can construct the desired $f_{t}$ by using unary functions of $\tilde{\Omega}/\tilde{\Omega}_{\rm SS}$, denoted by $\tilde{\psi}_{\rm c}$:
@@
In the derivation below, we additionally assume $\tilde{\psi}_{\rm c}$ to be a non-negative, monotonically decreasing function. Then, the monotonic increase of $x\ln x$ for $x\geq 1$, together with the monotonic decrease of $\tilde{\psi}_{\rm c}$, guarantees that eq. (115) holds during the transient evolution following positive steps, where $\tilde{\Omega}\geq\tilde{\Omega}_{\rm SS}$, as long as eq. (115) holds at steady states, where $\tilde{\Omega}=\tilde{\Omega}_{\rm SS}$. Hence, we only have to request $f_{t}$ of eq. (116) to conform eq. (115) during the transient evolution following negative steps:
$f_{t}=o\left[\frac{L}{D_{\rm c}}\frac{\tilde{\Omega}}{\tilde{\Omega}_{\rm SS}}\ln\frac{\tilde{\Omega}}{\tilde{\Omega}_{\rm SS}}\right]\hskip 10.0pt\left(\frac{\tilde{\Omega}}{\tilde{\Omega}_{\rm SS}}\gtrsim\frac{\min(V_{\rm after},V_{\rm X})}{\min(V_{\rm before},V_{\rm X})}\right)$ (117)
@@
Now we consider constructing a full evolution law as $f=f_{t}+f_{\delta}V$, using $f_{t}$ of eqs. (116) and $f_{\delta}V$ of eq. (112). Then, $f$ becomes a function of two variables $V$ and $\theta$. In the latter half of §4.2, we have shown that $f_{t}$ in such forms must satisfy eq. (86) for $\theta,\theta_{\rm SS}\gg\theta_{\rm X}$. Therefore, for $\tilde{\psi}_{\rm c}$ conforming eq. (118) to be usable in this construction, $f_{t}$ conforming eqs. (116) and (118) must also satisfy eq. (86) for $\theta,\theta_{\rm SS}\gg\theta_{\rm X}$. Recalling $\tilde{\Omega}/\tilde{\Omega}_{\rm SS}\simeq\Omega$ for $\theta,\theta_{\rm SS}\gg\theta_{\rm X}$, we see that such is possible only if
@@
Thus, we have shown that evolution laws including the effects of the strength minimum can be constructed as $f=f_{t}+f_{\delta}V$, using the $f_{\delta}$ of eq. (112) and the $f_{t}=\psi_{\rm c}(\tilde{\Omega}/\tilde{\Omega}_{\rm SS})$ with $\psi_{\rm c}$ of eq. (83), which satisfies eq. (118), a sufficient condition of eqs. (65) and (113):
@@
Notice that what fixes the $\tilde{\Omega}_{\rm SS}$ value is not eq. (121). The evolution law for $\theta$ is given as the combination of eqs. (111) and (121). In other words, transient evolution and value assignment to the steady state (as a function of $V$) are separable. This distinct characteristic is a direct consequence of part of the VS constraints, where $f_{\delta}$ (eq. 43) drives the system toward the steady state while $f_{t}$ disappears at the steady state, that is,
$f_{t}=o[(L/D_{\rm c})\tilde{\Omega}\ln(\tilde{\Omega}/\tilde{\Omega}_{\rm SS})]\hskip 10.0pt(\tilde{\Omega}\sim\tilde{\Omega}_{\rm SS}),$ (122)
which is a special case of eq. (113) in the vicinity of steady states.
We now check if the evolution law eq. (121) has a unique stationary point, one of the desired properties as discussed in §4.2. Because $\psi_{\rm c}$ is non-negative, the stationary points of eq. (121) can exist only in $\tilde{\Omega}/\tilde{\Omega}_{\rm SS}\geq 1$. Further recalling that $f_{t}$-part of eq. (121), $\psi_{\rm c}(\tilde{\Omega}/\tilde{\Omega}_{\rm SS})$ has been assumed to be monotonically decreasing, and also noticing that its $f_{\delta}V$-part, $-(\tilde{\Omega}/\tilde{\Omega}_{\rm SS})\ln(\tilde{\Omega}/\tilde{\Omega}_{\rm SS})$, is also monotonically decreasing in the range $\tilde{\Omega}/\tilde{\Omega}_{\rm SS}\geq 1$, we see that the trivial stationary point $\tilde{\Omega}/\tilde{\Omega}_{\rm SS}=1$, that is, $\Omega=\Omega_{\rm SS}$ is the only stationary point of eq. (121).
At the end of this subsection, we visually compare eq. (121) with eq. (80) derived in §4.3. Both are the sum of the slip-driven state evolution $f_{\delta}V$ and the purely time-dependent healing $f_{t}$ that disappears when strength comes to the close vicinity of steady states. We notice that the replacement of $\Omega$ in eq. (80) with $\tilde{\Omega}$ and another replacement of its steady-state value $C_{\rm SS}=1$ (eq. 76) with $\tilde{\Omega}_{\rm SS}$ (eq. 111) yields eq. (121). The former replacement corresponds to $\theta\to\tilde{\theta}$ and the latter corresponds to $L/V\to L/V+\theta_{\rm X}$. As expected from such relationships between them, it is easy to confirm that eq. (121) reproduces eq. (80) for $\theta,\theta_{\rm SS}\gg\theta_{\rm X}$, as it ought to. Thus, constructed eq. (121) (plus eq. 111) is modification of eq. (80) so that experimental requirements for $\theta\lesssim\theta_{\rm X}\cup\theta_{\rm SS}\lesssim\theta_{\rm X}$ are satisfied as well. In conclusion, the evolution law consisting of eqs. (83), (111), and (121) satisfies all the experimental requirements for the entire range of $\theta$ and $\theta_{\rm SS}$, including $\theta,\theta_{\rm SS}\gg\theta_{\rm X}$, $\theta\lesssim\theta_{\rm X}$, and $\theta_{\rm SS}\lesssim\theta_{\rm X}$, except the unavoidable side effects of aging-slip switch predicted outside the conditions experimentally probed so far.

### v3

@@ 4.5 Incorporation of the strength minimum
This state variable $\tilde{\theta}$ always has the same $\Phi$ dependency (eq. 16) as that of $\theta$ for the range $\theta\gg\theta_{\rm X}$ (eq. 12) mentioned above, and even for $\theta\ll\theta_{\rm X}$, the value of $\tilde{\theta}$ can be well constrained. Hence, we are able to derive meaningful evolution laws for $\tilde{\theta}$ from the canons. In fact, though counter-intuitive, the varying rate of $\theta$ (i.e., the evolution law in $\theta$ notation) is as well constrained as that of $\tilde{\theta}$ through $\dot{\theta}=\dot{\tilde{\theta}}$, despite the weak constraint on the value of $\theta$ for $\theta\ll\theta_{\rm X}$.
Now, we constrain the evolution law according to the canonical behaviors, this time considering the behaviors in the range $\theta\lesssim\theta_{\rm X}\cup\theta_{\rm SS}\lesssim\theta_{\rm X}$ as well. We begin with the SS canon, as we did so in the development for $\theta,\theta_{\rm SS}\gg\theta_{\rm X}$. For the entire range of $V$, including $V\gg L/\theta_{\rm X}$, $V\sim L/\theta_{\rm X}$, and $V\ll L/\theta_{\rm X}$, Weeks (1993) has shown that their experimental data of the steady-state shear stress $\tau_{\rm SS}$ are well fitted by eq. (10), namely the direct-effect term plus the steady-state strength, if the steady-state strength has the following velocity dependence involving a high-speed cutoff:
@@
This satisfies both the experimental requirement (eq. 26) for the SS test when $\theta_{\rm SS}\gg\theta_{\rm X}$ and the strength minimum (eq. 14). Through eq. (15), eq. (106) is equivalent to $\Omega_{\rm SS}=1$ (eq. 31). Therefore, considering the above result by Weeks (1993), we regard eq. (31) as the constraint from the SS canon for the entire range of $\theta$ and $\theta_{\rm SS}$.
The VS constraint (eq. 43) holds for any $\Phi$ and $\Phi_{\rm SS}$. Thus, no modification is necessary to incorporate the strength minimum. This eq. (43) is the only constraint on $f_{\delta}$, except that it must be negligible at the NSC, which is automatically met as long as eq. (43) on $f_{\delta}$ and eq. (65) on $f_{t|{\rm NSC}}$ hold. Hence, we can adopt the slip-dependent evolution function $f_{\delta{\rm|VS}}$ of eq. (43), which involves the effect of $\theta_{\rm X}$, as the evolution law $f_{\delta}$ satisfying the VS canon (eq. 38) for the entire range of $\theta$ and $\theta_{\rm SS}$.
For convenience in subsequent derivations, we introduce the following variable:
@@
Also, purely time-dependent healing under the NSC has already been constrained by eq. (65), which already considers the effect of $\theta_{\rm X}$ and is therefore valid for any $\theta$ and $\theta_{\rm SS}$:
@@
Thus, all that is missing is $f_{t}$ under the non-NSC. Similar to our derivation of eq. (83) for $\theta,\theta_{\rm SS}\gg\theta_{\rm X}$, the asymptotic form of $f_{t}$ under the non-NSC is constrained by eq. (44) of VS tests. With eq. (112), eq. (44) can be rewritten into
@@
When we derive eq. (112) from eq. (43), we have removed the conditional subscript ||||VS on the left-hand side of eq. (112) to indicate that eq. (112) represents the $f_{\delta}$ part of the evolution law being developed rather than the description of constraints in a specific type of experiment. Unlike eq. (112), eq. (113) is still an experimental constraint on the evolution law, so we keep ||||VS on the left-hand side of eq. (113). This is for the pragmatic reason of distinguishing it from another constraint but for the NSC (eq. 65).
@@
Thus, regarding the constraint on $f_{t}$, the NSC ($|f_{t}|\gg|f_{\delta}V|$) gives the range where the SHS canon applies while eq. (44) ($|f_{t}|\ll|f_{\delta}V|$) gives the range where the VS canon applies. On the other hand, regarding $f_{\delta}$, since the definition of the NSC ($|f_{t}|\gg|f_{\delta}V|$) means that $f_{\delta}$ is negligible, the explicit form of $f_{\delta}$ under the NSC is poorly constrained [from $\mathcal{O}(B)$ variations in $\Phi$]. However, this conversely means that we can assume eq. (112) for the entire range, including the NSC, without causing any relevant differences in the resultant behavior of $\Phi$. For this reason, in the present derivation, we can apply eq. (112) for the entire range. Similarly, noting that $f_{t}$ does not affect the evolution as long as $|f_{t}|\ll|f_{\delta}V|$ (eq. 44) holds, we can modify the constraint of eq. (113) to the following stricter one:
$f_{t|{\rm VS}}=o[(L/D_{\rm c})(\tilde{\Omega}/\tilde{\Omega}_{\rm SS})\ln(\tilde{\Omega}/\tilde{\Omega}_{\rm SS})],$ (114)
@@
where $f_{t}$ is more strongly muted than in eq. (113) because $\tilde{\Omega}_{\rm SS}\geq\Omega_{\rm SS}=1$ as seen from eqs. (31) and (111). Equation (114) is a sufficient condition of eq. (113), and the use of eq. (114) facilitates subsequent derivation.
@@
Using eq. (114), we can derive the evolution law by a logic parallel to that in §4.2. Specifically, by noting that $\tilde{\Omega}/\tilde{\Omega}_{\rm SS}$ connects to $\Omega$ when $\theta,\theta_{\rm SS}\gg\theta_{\rm X}$ (eqs. 30 and 105111), we can regard eq. (114) as a generalized form of $f_{t|{\rm VS}}=o[(L/D_{\rm c})\Omega\ln\Omega]$ derived from eqs. (44) and (72), where $\tilde{\Omega}/\tilde{\Omega}_{\rm SS}$ replaces $\Omega$ (i.e., $\Omega/\Omega_{\rm SS}$ given eq. 31). Therefore, $f_{t}$ satisfying eqs. (65) and (114) can be derived by repeating the same procedure as in §4.2, with $\Omega$ replaced by $\tilde{\Omega}/\tilde{\Omega}_{\rm SS}$.
@@
Since the right-hand side of eq. (114) is a unary function of $\tilde{\Omega}/\tilde{\Omega}_{\rm SS}$, eq. (114) is equivalent to that $f_{t}$ equals the right-hand side of eq. (114) in the range of $\tilde{\Omega}/\tilde{\Omega}_{\rm SS}$ that can occur in VS tests. In ideal VS tests, where the interface slip rate instantly changes from $V_{\rm before}$ to $V_{\rm after}$, the farthest deviation of the strength from the steady state occurs immediately after the step. Hence, we see that $|\ln(\tilde{\Omega}/\tilde{\Omega}_{\rm SS})|\leq|\ln[(L/V_{\rm before}+\theta_{\rm X})/(L/V_{\rm after}+\theta_{\rm X})|$ (eqs. 105, 110, and 111). This inequality obtained for ideal VS tests is approximately valid for realistic (non-ideal) VS tests. Thus, eq. (114) can be rewritten into
@@
$f_{t}=o\left[\frac{L}{D_{\rm c}}\frac{\tilde{\Omega}}{\tilde{\Omega}_{\rm SS}}\ln\frac{\tilde{\Omega}}{\tilde{\Omega}_{\rm SS}}\right]\hskip 10.0pt\left(\left(\ln\frac{\min(V_{\rm after},V_{\rm X})}{\min(V_{\rm before},V_{\rm X})}\right)\left/\left(\ln\frac{\tilde{\Omega}}{\tilde{\Omega}_{\rm SS}}\right)\gtrsim\right.1\right),$ (115)
@@
Lastly, we look for $f_{t}$ satisfying both eq. (65) (constraint for the NSC) and eq. (115) (constraint for $|f_{t}|\ll|f_{\delta}V|$). Since the right-hand side of eq. (115) is a unary function of $\tilde{\Omega}/\tilde{\Omega}_{\rm SS}$, we can construct the desired $f_{t}$ by using unary functions of $\tilde{\Omega}/\tilde{\Omega}_{\rm SS}$, denoted by $\tilde{\psi}_{\rm c}$:
@@
In the derivation below, we additionally assume $\tilde{\psi}_{\rm c}$ to be a non-negative, monotonically decreasing function. Then, the monotonic increase of $x\ln x$ for $x\geq 1$, taken together with the monotonic decrease of $\tilde{\psi}_{\rm c}$, guarantees that eq. (115) holds during the transient evolution following positive steps, where $\tilde{\Omega}\geq\tilde{\Omega}_{\rm SS}$, as long as eq. (115) holds at steady states, where $\tilde{\Omega}=\tilde{\Omega}_{\rm SS}$. Hence, we only have to request $f_{t}$ of eq. (116) to conform eq. (115) during the transient evolution following negative steps:
$f_{t}=o\left[\frac{L}{D_{\rm c}}\frac{\tilde{\Omega}}{\tilde{\Omega}_{\rm SS}}\ln\frac{\tilde{\Omega}}{\tilde{\Omega}_{\rm SS}}\right]\hskip 10.0pt\left(\frac{\tilde{\Omega}}{\tilde{\Omega}_{\rm SS}}\gtrsim\frac{\min(V_{\rm after},V_{\rm X})}{\min(V_{\rm before},V_{\rm X})}\right).$ (117)
@@
Now, we consider constructing a full evolution law as $f=f_{t}+f_{\delta}V$, using $f_{t}$ of eqs. (116) and (118) and $f_{\delta}V$ of eq. (112). Then, $f$ becomes a function of two variables $V$ and $\theta$. In the latter half of §4.2, we have shown that $f_{t}$ in such forms must satisfy eq. (86) for $\theta,\theta_{\rm SS}\gg\theta_{\rm X}$. Therefore, for $\tilde{\psi}_{\rm c}$ conforming eq. (118) to be usable in this construction, $f_{t}$ conforming eqs. (116) and (118) must also satisfy eq. (86) for $\theta,\theta_{\rm SS}\gg\theta_{\rm X}$. Recalling $\tilde{\Omega}/\tilde{\Omega}_{\rm SS}\simeq\Omega$ for $\theta,\theta_{\rm SS}\gg\theta_{\rm X}$, we see that such is possible only if
@@
Thus, we have shown that evolution laws incorporating the strength minimum can be constructed as $f=f_{t}+f_{\delta}V$, using the $f_{\delta}$ of eq. (112) and the $f_{t}=\psi_{\rm c}(\tilde{\Omega}/\tilde{\Omega}_{\rm SS})$ with $\psi_{\rm c}$ of eq. (83), which satisfies eq. (118), a sufficient condition of eqs. (65) and (113):
@@
Notice that what fixes the $\tilde{\Omega}_{\rm SS}$ value is eq. (111), not eq. (121). The evolution law for $\theta$ is given as the combination of eqs. (111) and (121). In other words, transient evolution and value assignment to the steady state (as a function of $V$) are separable. This distinct characteristic is a direct consequence of part of the VS constraints, where $f_{\delta}$ (eq. 43) drives the system toward the steady state while $f_{t}$ disappears at the steady state, that is, $f_{t}=o[(L/D_{\rm c})\tilde{\Omega}\ln(\tilde{\Omega}/\tilde{\Omega}_{\rm SS})]$ for $(\tilde{\Omega}\sim\tilde{\Omega}_{\rm SS}),$which is a special case of eq. (113) in the vicinity of steady states.
We now check if the evolution law eq. (121) has a unique stationary point, one of the desired properties as discussed in §4.2. Because $\psi_{\rm c}$ is non-negative, the stationary points of eq. (121) can exist only in $\tilde{\Omega}/\tilde{\Omega}_{\rm SS}\geq 1$. Further recalling that $f_{t}$-part of eq. (121), $\psi_{\rm c}(\tilde{\Omega}/\tilde{\Omega}_{\rm SS})$, has been assumed to be monotonically decreasing, and also noticing that its $f_{\delta}V$-part, $-(\tilde{\Omega}/\tilde{\Omega}_{\rm SS})\ln(\tilde{\Omega}/\tilde{\Omega}_{\rm SS})$, is also monotonically decreasing in the range $\tilde{\Omega}/\tilde{\Omega}_{\rm SS}\geq 1$, we see that the trivial stationary point $\tilde{\Omega}/\tilde{\Omega}_{\rm SS}=1$, that is, $\Omega=\Omega_{\rm SS}$, is the only stationary point of eq. (121).
At the end of this subsection, we visually compare eq. (121) with eq. (80) derived in §4.3. Both are the sum of the slip-driven state evolution $f_{\delta}V$ and the purely time-dependent healing $f_{t}$ that disappears when strength comes to the close vicinity of steady states. We notice that the replacement of $\Omega$ in eq. (80) with $\tilde{\Omega}$ and another replacement of its steady-state value $C_{\rm SS}=1$ (eq. 76) in $\ln\Omega(=\ln(\Omega/C_{\rm SS}))$ with $\tilde{\Omega}_{\rm SS}$ (eq. 111) yields eq. (121). The former replacement corresponds to $\theta\to\tilde{\theta}$, and the latter corresponds to $L/V\to L/V+\theta_{\rm X}$. As expected from such relationships between them, it is easy to confirm that eq. (121) reproduces eq. (80) for $\theta,\theta_{\rm SS}\gg\theta_{\rm X}$, as it ought to. Thus, the constructed eq. (121) (plus eq. 111) is the modification of eq. (80) so that experimental requirements for $\theta\lesssim\theta_{\rm X}\cup\theta_{\rm SS}\lesssim\theta_{\rm X}$ are satisfied as well. In conclusion, the evolution law consisting of eqs. (83), (111), and (121) satisfies all the experimental requirements for the entire range of $\theta$ and $\theta_{\rm SS}$, including $\theta,\theta_{\rm SS}\gg\theta_{\rm X}$, $\theta\lesssim\theta_{\rm X}$, and $\theta_{\rm SS}\lesssim\theta_{\rm X}$, except the unavoidable side effects of aging-slip switch predicted outside the conditions experimentally probed so far.